OpenStudy (anonymous):

h is to 8 as 3 is to 12. What is h? Help with this proportion please it's pretty simple but I drew a blank:P please help, thanks:D

OpenStudy (phi):

h is to 8 as 3 is to 12 write "h is to 8" as a fraction h/8 write 3 is to 12 as a fraction 3/12 \[ \frac{h}{8}= \frac{3}{12} \] you should notice that 3 divides into both 3 (itself) and 12, so you can simplify 3/12 divide top by 3 to get 1, divide bottom 12 by 3 to get 4 \[ \frac{h}{8}= \frac{1}{4} \] to solve for h, multiply both sides by 8 and simplify: \[ \frac{h}{8}\cdot 8 = \frac{1}{4} \cdot 8\]

OpenStudy (anonymous):

It is ok now I got the answer, it is 2, but thank you! :D I got it by writing how you said 3/12 and h/8 and then I cross-multiplied and got 12xH and 8x3 then I divided by 12 too both sides and got 12h/12 = 24/12 and the answer was H = 2 Thank You!!!:D